UARS Weekly Status Report

08 February 2002




********************************************************************

                UARS WEEKLY STATUS REPORT - 08 Feb 2002

********************************************************************

***********
SOLSTICE
***********

With the exception of Day only operations during the yaw maneuver this
week, SOLSTICE continues to operate normally, taking solar data on all
available orbital tracking periods.

***********
SUSIM
***********

On 2/2/02, SUSIM was turned off, as usual, to conserve power during the
UARS yaw-around maneuver.  On 2/6/02, SUSIM was repowered and on
2/7/02, daily mid-resolution and MgII solar scans resumed scheduled
during periods of predicted UARS telemetry.

Routine closed loop SSPP pointing was begun on Jan. 30.  Preliminarily,
the results are terrific (< 1 arcmin) as expected.  Of course, this
does not in any way diminish the miraculous improvement in the open
loop pointing and attitude control provided earlier by the UARS Project
Engineers.  Great job!

Although current and some previous level 0 data have now been delivered
from the PACOR-A system, the backlog of level 0 data from the last few
months has not arrived, for the most part.

After several missing features and incompatibilities were overcome,
the software conversion of SUSIM's STOL procedures to the CECIL
system is complete and is awaiting testing.

***********
ACRIM
***********

Nothing received.

***********
PEM
***********

     A detailed review of the PEM CICEL PROCs were conducted.  10 of 51 PROCs 
were found to be missing.  These need to be included in the CECIL conversion.
Of the 41 remaining PROCs, 39 of the CECIL PROCs were modified.  PEM has 
resubmitted the 39 CECIL PROCs and is waiting for the 10 missing CECIL PROCs 
to be sent to PEM for review.  PEM would like to know if the instrument 
activation STOL procedures archived in the FOT are to be converted by GSFC.  

     PEM continues to receive NMC data files from NOAA.  No problems have
been encountered.

     PEM awaits support of the UARS Traceability Mission for the TIMED mission
and has requested meetings with a GUVI team representative.

*Beginning of power orbit day (PEM activation): array current >5 amps, array 
     voltage >53.5 volts, and battery current >2.0 amps.  
 End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
     array current <9 amps, and battery currents <0.2 amps.
 Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.

***********
HALOE 
***********

Nothing received.

***********
MLS
***********

Nothing received.

***********
HRDI
***********

Nothing received.

***********
WINDII
***********

Nothing received.

***********
MPG
***********

The MPG participated in the UARS Ground System
Rehost Meeting on 7 February 2002.

The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop 
capabilities for the Science Traceability Mission.

Daily processing of instrument Daily Activity Plans 
continues with no problems.

***********
CDHF
***********

Nothing received.
                  
***********
FOT
***********

                           REPORTING PERIOD

This  report covers 26 January 2002 (Orbit 56733, GMT Day 026) through
01 February 2002 (Orbit 56837, GMT Day 032).

                         SPACECRAFT OPERATION

The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and   is  performing  nominally  under  two-battery  operations.   The
instrument operational changes were:


                    INSTRUMENT OPERATIONAL CHANGES


       Instrument    Time          Orbit     Comment
       PEM           026/2323      56747          AXIS-1 Off
       HALOE         031/2323      56822          Power Off


The  beta  angle ranged between the angles of 34.53 degrees  to  09.65
degrees  for  this  report period.  The beta angle is  now  decreasing
towards a minimum angle of 0 degrees on 03 February 2002 (DOY 034).
   
Spacecraft  battery  1,  2,  and 3 performance  monitoring  continues.
Battery 1 remains OFF the charge relay.  Battery 2 and Battery  3  are
currently maintaining greater than 25.44 V end-of-night (EON) load bus
voltage.   Battery  temperatures are stable with a  temperature  delta
between Battery 2 and Battery 3 of 1.31 to 1.25 degrees C.  Additional
battery  performance data for each battery on 26 January 2002  and  01
February 2002 are provided in Appendix D.

The  clock  error  ranged from -11.4 to 6.1 msec  during  this  report
period.  Clock rate adjustments are listed in Appendix C.

All  instrument  and subsystem engineers were advised on  the  initial
excursion  of ALERTS or Out-Of-Limits conditions detected by  the  FOT
for  this report.  A summary is listed below and the details of  ALERT
and Out-Of-Limits occurrences are listed in Appendix B.

    OUT-OF-LIMITS  =    00
    ALERTS         =    00

The  following real-time command sheets were executed for  the  listed
operational element(s) during this report period (see Appendix A).

    FOT            =    1
    PEM            =    2
    HALOE          =    1

UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.

    FOT            =    0

Zero  (0)  Anomaly  Reports remained OPEN at the end  of  this  report
period (see Appendix E).

TRANSPONDER FREQUENCIES:
    Transponder A:  01/333  - 2287.496855
    Transponder B:  01/324  - 2287.498688


                        GROUND SYSTEM OPERATION


Zero  (0) operational support problems occurred this period.   Anomaly
Reports are included in Appendix E.

The  data  loss  calculations are no longer  being  supplied  by  Data
Capture Facility (DCF).



                       OTHER SIGNIFICANT EVENTS


                       UARS SATELLITE OPERATIONS


The FOT tested loading capability of Eclipse versus the simulator this
week.   The FOT was able to successfully perform a partial table  load
without  checksum, spacecraft and TDRS ephemeris loads, ATC loads,  an
RTS  load  and a SUSIM microprocessor load.  This was a great  success
for Eclipse and keeps the reengineer effort on schedule for spacecraft
loads starting on 06 February 2002.  The first load to be sent to UARS
will be a partial table load.


                               MEETINGS

The  FOT  attended  a  UARS Ground System Rehost  Meeting  on  Monday,
January 28, 2002.

The  FOT  attended a UARS Eclipse Load Directory Structure  Review  on
Wednesday, January 30, 2002.

The  FOT  attended a UARS Eclipse Working Meeting on Thursday, January
31, 2002.


                       FUTURE SIGNIFICANT EVENTS


The  next  UARS Yaw slew is scheduled for 03 February 2002 (DOY  034).
This  maneuver will be a Reverse to Forward Flight maneuver  (Southern
Hemisphere viewing).


Attachments:    Appendix A, Real-time Command Sheets
                Appendix B, Out-Of-Limits & ALERT Occurrences
                Appendix C, Clock & Solar Array Rate Corrections
                Appendix D, UARS Battery Performance Data
                Appendix E, UARS Anomaly Reports
                              APPENDIX A
                       REAL-TIME COMMAND SHEETS
                     Week ending 01 February 2002

ORBIT   AOS        NO.      SUBSYS      REASON
-----   ---        ---      ------      ------
56747   026/2323   02-027   PEM         AXIS-1 Off
56822   031/2323   02-028   OBC/CLK     Clk Adj from 0.00 to -20.6 msec/day
56822   031/2323   02-029   HALOE       Power Off
56837   032/2247   02-030   PEM         Power Off
                                   
                              APPENDIX B
                  OUT-OF-LIMITS & ALERTS OCCURRENCES
                     Week ending 01 February 2002

The following Out-Of-Limits occurrences were detected:

MNEMONIC       STATE   DESCRIPTION                ORBITS
--------       -----   -----------                ------
-None this report period.


The following ALERT occurrences were detected:

MNEMONIC     DESCRIPTION                          ORBITS
--------     -----------                          ------
-None this report period.

______________________________________________________________________

                              APPENDIX C
                 CLOCK & SOLAR ARRAY RATE CORRECTIONS
                     Week ending 01 February 2002

                     CLOCK DRIFT RATE ADJUSTMENTS

ORBIT          TIME                   CLOCK RATE CHANGE
-----          -----                  -----------------
56822          031/2323               Clk Adj from 0.00 to -20.6 msec/day


                              APPENDIX D
                     UARS BATTERY PERFORMANCE DATA
                     Week ending 01 February 2002


  26 January 2002 - GMT Day 026 Beta = 34.53 deg, SA Pos = 269 deg

     DIFF V (mV) TEMP (C)     CURR (amp)  EON V    AVE C/D  AVE
BAT  MAX/MIN     MAX/MIN      MAX/MIN     MIN      RATIO    DOD(%)
---  -------     --------     ----------  -----    -------  ------
1  -50.4/-67.2  -3.77/-4.69  +0.0/ -0.0   10.72    0.000    0.00
2  +11.2/-33.6  +5.14/+4.48  +28.0/-14.4  25.44    1.020    20.97
3  +28.0/-16.8  +4.15/+3.17  +27.2/-14.0  25.44    1.020    20.62


  01 February 2002 - GMT Day 032 Beta = 09.65 deg, SA Pos = 269 deg


     DIFF V (mV) TEMP (C)     CURR (amp)  EON V    AVE C/D  AVE
BAT  MAX/MIN     MAX/MIN      MAX/MIN     MIN      RATIO    DOD(%)
---  -------     ----------   ----------  -----    -------  ------
1  -44.8/-67.2  -4.08/-4.99  +0.0/ -0.0   10.72    0.000    0.00
2  +0.0/-33.6   +3.82/+3.50  +28.4/-12.8  26.40    1.030    19.22
3  +22.4/-22.4  +2.85/+2.20  +27.2/-12.4  26.24    1.030    18.64

      V/T LEVEL CHANGES
ORB#          TIME       LVL FR-TO
-----         ----       ---------
-no changes    V/T 5

   





                              APPENDIX E
                         UARS ANOMALY REPORTS
                     Week ending 01 February 2002

                     NEW ANOMALY REPORTS GENERATED
AR NO.    SUBSYS    ORBIT     TIME      TTR       PROBLEM/DATA LOSS
------    ------    -----     ----      ---       -----------------
-None this report period.

                ANOMALY CLOSURE INFORMATION RECEIVED
AR NO.    SUBSYS    ORBIT     TIME      TTR       PROBLEM/DATA LOSS
------    ------    -----     ----      ---       -----------------
02-002    FOT       see list  Jan 02    14170     Generic Late Acquisition

                  ANOMALY INVESTIGATION REPORTS GENERATED
AR NO.    SUBSYS    ORBIT     TIME      TTR       PROBLEM/DATA LOSS
------    ------    -----     ----      ---       -----------------
-None this report period.

                         OPEN ANOMALY REPORTS
AR NO.    SUBSYS    ORBIT     TIME      TTR       PROBLEM/DATA LOSS
------    ------    -----     ----      ---       -----------------
-None this report period.